# UniST: A Prompt-Empowered Universal Model for Urban Spatio-Temporal Prediction

### v3

@@ Abstract.
Urban spatio-temporal prediction is crucial for informed decision-making, such as transportation management, resource optimization, and urban planning. Although pretrained foundation models for natural languages have experienced remarkable breakthroughs, wherein one general-purpose model can tackle multiple tasks across various domains, urban spatio-temporal modeling lags behind. Existing approaches for urban prediction are usually tailored for specific spatio-temporal scenarios, requiring task-specific model designs and extensive in-domain training data. In this work, we propose a universal model, UniST, for urban spatio-temporal prediction. Drawing inspiration from large language models, UniST achieves success through: (i) flexibility towards diverse spatio-temporal data characteristics, (ii) effective generative pre-training with elaborated masking strategies to capture complex spatio-temporal relationships, (iii) spatio-temporal knowledge-guided prompts that align and leverage intrinsic and shared knowledge across scenarios. These designs together unlock the potential of a one-for-all model for spatio-temporal prediction with powerful generalization capability. Extensive experiments on 15 cities and 6 domains demonstrate the universality of UniST in advancing state-of-the-art prediction performance, especially in few-shot and zero-shot scenarios. The implementation is available at this repository: https://github.com/tsinghua-fib-lab/UniST.
Figure 1. From traditional separate deep models to a one-for-all universal model for urban spatio-temporal prediction.
@@ 1. Introduction
Pre-trained foundation models have showcased remarkable success in Natural Language Processing (NLP) (Touvron et al., 2023, 2023; Brown et al., 2020), particularly excelling in few-shot and zero-shot scenarios (Brown et al., 2020; Kojima et al., 2022). However, the field of urban spatio-temporal modeling (Wang et al., 2020; Zheng et al., 2014; Jin et al., 2023b) has not witnessed a parallel breakthrough. In this paper, our goal is to build a foundation model for urban spatio-temporal prediction, that is, to develop a universal spatio-temporal model, i.e., a one-for-all model, with superior performance and robust generalization capability across diverse spatio-temporal scenarios. These scenarios encompass a wide range of domains, such as transportation, cellular network, and human mobility, spanning various cities.
@@
The significance of such a universal model lies in its ability to address prevalent data scarcity issues in urban areas, where varying degrees of digitization result in imbalanced data collections across different cities and domains. While impressive progress has been made in existing spatio-temporal modeling approaches (Bai et al., 2020; Zhang et al., 2017; Liu et al., 2018; Cheng et al., 2023; Bai et al., 2019a; Zhou et al., 2023b; Xia et al., 2023; Pan et al., 2019; Zhou et al., 2023d), their effectiveness is often constrained by rigid specialization to a particular domain in one city, i.e., a specific dataset. The reliance on extensive training data further impedes the model’s generalization potential. Consequently, current solutions are still far from “universality”, and have restrictive applications.
@@
A universal spatio-temporal model should have the following two capabilities. Firstly, it must be capable of leveraging abundant and rich data from urban scenarios for training, particularly data characterized by varying spatio-temporal features. The training of the foundational model should ensure the acquisition of ample and rich information (Touvron et al., 2023; Wang et al., 2023; Bai et al., 2023). Second, it should demonstrate robust generalization across diverse spatio-temporal scenarios. Especially in scenarios with limited or no training data, the model can still predict well without much performance degradation (Garza and Mergenthaler-Canseco, 2023; Kirillov et al., 2023; Wang et al., 2023). Crucially, this capability emerges not only from diverse data training, but also from the shared and transferable knowledge across scenarios.
@@
However, realizing the aforementioned two characteristics is technically non-trivial. Several challenges unique to spatio-temporal data prevent the direct adoption of current foundation models from the language or vision domains. The first challenge arises from diverse data formats inherent in spatio-temporal datasets. Unlike languages with a natural and unified 1D structure or images and videos adhering to standard formats, spatio-temporal data collected from different sources exhibit completely different features. The variable dimension, temporal duration, and spatial range may vary significantly, making it difficult to impose a standardized structure. The second challenge arises from the high variations in data distributions across multiple scenarios. Faced with highly distinct spatio-temporal patterns, the model may struggle to adapt to these differences. Moreover, unlike NLP, which benefits from a shared vocabulary, different spatio-temporal scenarios often have entirely different spatial and temporal scales, lacking common elements for knowledge sharing and transfer.
Although the displayed spatio-temporal patterns vary significantly, we believe that some underlying principles are shared among these diverse scenarios. For instance, the locations of commercial centers in different cities are probably different, yet the way they correlate with other functional districts could be a shared characteristic. Similarly, while temporal periodic patterns differ across domains, they share the essential concept of repetitive patterns. Therefore, the key point of building a one-for-all model is to align and leverage these shared but underlying characteristics.
We propose UniST, a universal solution for urban spatio-temporal prediction. Notably, it ingeniously imitates the key features of LLMs:
@@
    Flexibility towards diverse data characteristics;
@@
    Effective pre-training to capture multi-level spatio-temporal relationships;
@@
    Spatio-temporal knowledge-guided prompt to align underlying and shared commonalities across scenarios.
UniST achieves the above characteristics through its holistic design, which is specified by four key components: data, architecture, pre-training, and prompt tuning. Firstly, we seek to harness the rich diversity inherent in spatio-temporal scenarios by collecting extensive data from various cities and domains. Secondly, we employ spatio-temporal tokenizers to convert diverse data shapes into a unified sequential format, which facilities to utilize the powerful Transformer architecture. Thirdly, drawing inspiration from language (Devlin et al., 2018; Brown et al., 2020) and vision communities (He et al., 2022), UniST adopts the widely-used generative pre-training strategy – masked token modeling. Going beyond, we employ different masking strategies tailored for spatio-temporal data, refining the model’s ability to comprehensively capture spatio-temporal correlations. Moreover, we introduce an innovative spatio-temporal knowledge-guided prompt mechanism. Inspired by well-established domain knowledge in spatio-temporal modeling, an elaborated prompt network identifies underlying and shared spatio-temporal properties, and learns to adaptively leverage them to generate useful prompts. In this way, we can align distinct data distributions of various datasets and develop a one-for-all universal model. We summarize our contributions as follows:
  • To the best of our knowledge, this is the first attempt to address the universal spatio-temporal prediction, which investigates the potential of a one-for-all model in diverse urban scenarios.
  • We propose UniST, a universal spatio-temporal prediction model. Through the elaborated masking strategies and spatio-temporal knowledge-guided prompts, it manages to harness the data diversity and achieve a universal solution for all scenarios. It represents a paradigm shift in spatio-temporal prediction from traditional separate methods to pre-trained foundational models.
  • Extensive experiments demonstrate the strong universality of UniST. It achieves new state-of-the-art performance on various prediction tasks. Particularly, it demonstrates superior transferability to previously unseen scenarios.
@@ 2. Related Work
Urban spatio-temporal prediction (Wang et al., 2020; Zheng et al., 2014; Jin et al., 2023b) aims to model and forecast the dynamic patterns of urban activities over both space and time. The advent of deep learning techniques has propelled significant advancements in urban spatio-temporal prediction. A spectrum of models, including CNN-based (Li et al., 2018; Zhang et al., 2017; Liu et al., 2018), RNN-based (Wang et al., 2017, 2018a; Lin et al., 2020; Bai et al., 2019c), ResNet-based (Zhang et al., 2017), GNN-based (Zhao et al., 2019; Jin et al., 2023b; Bai et al., 2020, 2019a, 2019b; Geng et al., 2019), Transformer-based (Chen et al., 2022; Jiang et al., 2023a; Yu et al., 2020; Chen et al., 2021), MLP-based (Zhang et al., 2023c; Shao et al., 2022) and diffusion model-based (Zhou et al., 2023a; Wen et al., 2023; Yuan et al., 2023) architectures, has been proposed to capture spatio-temporal relationships. Simultaneously, cutting-edge learning techniques like meta-learning (Chen et al., 2023; Lu et al., 2022; Yao et al., 2019), contrastive learning (Ji et al., 2023; Zhang et al., 2023b; Tang et al., 2023), and adversarial learning (Ouyang et al., 2023; Gao et al., 2022b; Fang et al., 2022; Tang et al., 2022) are also utilized. However, most existing approaches remain constrained by training separate models for each specific domain. Some studies (Yao et al., 2019; Lu et al., 2022; Jin et al., 2022; Wang et al., 2018b; Jin et al., 2023a; Jiang et al., 2023b; Liu et al., 2023e; Chen et al., 2023; Fang et al., 2022) explore transfer learning between cities, however, a certain amount of data samples in the target city are still required. Current solutions are restrictive to specified domains or cities that have training data. On the contrary, our model allows training and predicting across diverse spatio-temporal scenarios and provides a universal solution.
@@
Inspired by the remarkable strides in foundation models for NLP (Touvron et al., 2023, 2023; Brown et al., 2020) and CV (Rombach et al., 2022; Bai et al., 2023), the exploration of foundation models for urban scenarios has emerged as a compelling problem. Current explorations mainly unlock the potential of LLMs in this context. State-of-the-art urban systems like TransGPT (Tra, 2023), CityGPT (Xu et al., 2023), and TrafficGPT (Zhang et al., 2023a; Da et al., 2023) have demonstrated proficiency in addressing language-based spatio-temporal tasks. Additionally, LLMs are utilized to articulate descriptions for urban-related images (Yan et al., 2023), which is useful for downstream tasks. Moreover, the application of LLMs extends to traffic signal control (Lai et al., 2023), showcasing their utility in tackling complex spatio-temporal problems beyond languages. Recently, there has been great progress in foundation models for time series modeling (Jin et al., 2023d; Cao et al., 2023; Jin et al., 2023c; Liu et al., 2023a; Zhou et al., 2023c). Unlike time series characterized by a straightforward 1D structure, spatio-temporal data presents a more intricate nature with intertwined dependencies across both spatial and temporal dimensions. We think that spatio-temporal data is not inherently reliant on language. While the integration of LLMs is both interesting and worthy of exploration, developing a foundation model that specifically goes from spatio-temporal data, also holds significant importance.
Prompt Tuning Techniques
@@
Prompting techniques have achieved superior performance in Natural Language Processing (NLP) (Liu et al., 2023d; Qiao et al., 2022) and Computer Vision (CV) (Jia et al., 2022; Shen et al., 2024) fields. The goal is to enhance the generalization capability of pretrained models on specific tasks or domains. Typically, language models usually use a limited number of demonstrations as prompts and vision models often employ a learnable prompt network to generate useful prompts, known as prompt tuning. Our research aligns with prompt tuning, where spatio-temporal prompts are adaptively generated based on the spatio-temporal properties through a prompt network.
3. Preliminary
@@
Figure 2. The overview of UniST, which consists of two stages: (i) large-scale spatio-temporal pre-trianing, (ii) spatio-temporal knowledge-guided prompt tuning.
Spatial Partition and Temporal Aggregation.
The spatial partition and temporal aggregation vary across different spatio-temporal scenarios. In this study, we focus on grid-based spatial partitioning, where the spatial dimension is discretized into non-overlapping grids of equal size based on longitude and latitude. The grid map takes the form of an $H\times W$ rectangle, with each grid representing a distinct geographical region within the city. Temporal aggregation involves collecting data over specific temporal intervals. For example, crowd flows are documented on an hourly basis, while bike usage data is logged every half an hour.
Spatio-temporal Data.
With the spatial partition and temporal aggregation described above, spatio-temporal data in our study can be defined as a four-dimensional tensor: $T\times C\times H\times W$, where $T$ represents the number of time periods, $H$ and $W$ represent the number of grids in the longitude and latitude dimensions of the spatial partition, respectively, and $C$ represents the number of variables. Notably, $T$, $C$, $H$, and $W$ often vary across scenarios due to different sensors and implementations driven by specific objectives.
Spatio-temporal Prediction.
For a specific dataset, given $L_{H}$ historical observations $X$ for the grid map, we aim to predict the future $k$ steps. The spatio-temporal prediction task can be formulated as learning a $\theta$-parameterized model $\mathcal{F}$: $X_{[t:t+k]}=\mathcal{F}_{\theta}(X_{[t-L_{H}:t]})$.
Spatio-temporal Few-shot Prediction.
Consider a collection of source datasets denoted as ${X^{1},X^{2},...,X^{N}}$, and a target dataset $X^{tar}$ characterized by limited data. After training a model on the source datasets, the model can leverage acquired knowledge to adapt to the target dataset, even with limited training data.
Spatio-temproal Zero-shot Prediction.
The model can directly predict without additional training on the target dataset.
@@ 4. The UniST Model
We propose the UniST model to address universal spatio-temporal prediction, which is shown in Figure 2. It consists of two stages:
  • Stage 1: Large-scale spatio-temporal pre-training. Different from existing methods limited to a single dataset, our approach involves collecting extensive spatio-temporal data from a variety of cities and domains for pre-training. Concurrently, we introduce innovative spatio-temporal masking and reconstruction strategies to further enhance the pre-training process.
  • Stage 2: Spatio-temporal knowledge-guided prompt tuning. This stage introduces a prompt network for in-context learning. The generation of prompts is adaptively guided by well-developed spatio-temporal domain knowledge. We fix the feed-forward and attention layers of the base Transformer model. Subsequently, we fine-tune both the prompt network and the remaining parameters of the Transformer model.

### v4

@@ Abstract.
Urban spatio-temporal prediction is crucial for informed decision-making, such as traffic management, resource optimization, and emergence response. Despite remarkable breakthroughs in pretrained natural language models that enable one model to handle diverse tasks, a universal solution for spatio-temporal prediction remains challenging. Existing prediction approaches are typically tailored for specific spatio-temporal scenarios, requiring task-specific model designs and extensive domain-specific training data. In this study, we introduce UniST, a universal model designed for general urban spatio-temporal prediction across a wide range of scenarios. Inspired by large language models, UniST achieves success through: (i) utilizing diverse spatio-temporal data from different scenarios, (ii) effective pre-training to capture complex spatio-temporal dynamics, (iii) knowledge-guided prompts to enhance generalization capabilities. These designs together unlock the potential of building a universal model for various scenarios. Extensive experiments on more than 20 spatio-temporal scenarios demonstrate UniST’s efficacy in advancing state-of-the-art performance, especially in few-shot and zero-shot prediction. The datasets and code implementation are released on https://github.com/tsinghua-fib-lab/UniST.
Spatio-temporal prediction, prompt learning, universal model
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671662isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
Pre-trained foundation models have showcased remarkable success in Natural Language Processing (NLP) (Touvron et al., 2023; Brown et al., 2020), particularly excelling in few-shot and zero-shot settings (Brown et al., 2020; Kojima et al., 2022). However, similar breakthroughs have not yet been achieved in the field of urban spatio-temporal prediction (Wang et al., 2020; Zheng et al., 2014; Gong et al., 2023). In this paper, our goal is to establish a foundation model for general urban spatio-temporal prediction — specifically, to develop a universal model that offers superior performance and powerful generalization capabilities across diverse spatio-temporal scenarios. This entails training a single model capable of effectively handling various urban contexts, encompassing various domains such as human mobility, traffic and communication networks across different cities.
@@
The significance of such a universal model lies in its ability to address prevalent data scarcity issues in urban areas. The varying levels of digitalization across domains and cities often result in imbalanced and incomplete datasets. Despite notable advancements in existing spatio-temporal modeling approaches (Bai et al., 2020; Zhang et al., 2017; Liu et al., 2018; Pan et al., 2019; Zhou et al., 2023c; Li et al., 2023a), their effectiveness is typically confined to specific domains within a single city. The reliance on extensive training data further impedes the model’s generalization potential. Consequently, current solutions are still far from “universality”, and remain narrowly applicable.
@@
A universal spatio-temporal model must possess two essential capabilities. Firstly, it must be capable of leveraging abundant and rich data from different urban scenarios for training. The training of the foundational model should ensure the acquisition of ample and rich information (Touvron et al., 2023; Wang et al., 2023a; Bai et al., 2023). Second, it should demonstrate robust generalization across different spatio-temporal scenarios. Especially in scenarios with limited or no training data, the model can still work well without obvious performance degradation (Garza and Mergenthaler-Canseco, 2023; Wang et al., 2023a).
Figure 1. The transition from traditional separate deep learning models to a one-for-all universal model for urban spatio-temporal prediction.
@@
However, realizing the aforementioned capabilities encounters significant challenges specific to spatio-temporal data, which impede the direct application of current foundation models developed for language and vision domains. The first challenge arises from the inherent diverse formats of spatio-temporal datasets. Unlike languages with a natural and unified sequential structure or images and videos adhering to standardized dimensions, spatio-temporal data collected from different sources exhibit highly varied features. These include variable dimensions, temporal durations, and spatial coverages that differ significantly, posing difficulties in standardizing their structure. The second challenge arises from high variations in data distributions across multiple scenarios. Faced with highly distinct spatio-temporal patterns, the model may struggle to adapt to these differences. Unlike language, which benefits from a shared vocabulary, various scenarios of different domains and cities often operate on entirely different spatial and temporal scales, lacking common elements for effective training and generalization.
Although the displayed spatio-temporal patterns vary significantly, there are certain underlying laws that should be common among them. This principle arises from the intuition that human activity influences various spatio-temporal data generated in urban settings, leading to the existence of universal patterns. For example, traffic speed and communication networks exhibit distinct spatio-temporal patterns, yet both are influenced by human mobility and therefore adhere to similar underlying principles. Additionally, while temporal periodic patterns vary across domains, they share fundamental concept of repetition. Furthermore, city layouts vary considerably between different urban areas, but the relationships among various functional zones within cities may exhibit shared characteristics. Therefore, the key to building a one-for-all model is to capture, align and leverage these shared while underlying characteristics effectively.
To this end, we introduce UniST, a universal solution for urban spatio-temporal prediction through advanced pre-training and prompt learning. Notably, UniST achieves three essential capabilities of:
@@
    Scalability across scenarios with diverse spatio-temporal data;
@@
    Effective pre-training to capture complex spatio-temporal relationships;
@@
    utilizing spatio-temporal prompts to align underlying shared patterns across scenarios.
UniST achieves the above capabilities through its holistic design driven by four key components: data, architecture, pre-training, and prompt learning. Firstly, we harness the rich diversity inherent in spatio-temporal scenarios by leveraging extensive data from various domains and cities. Secondly, we design spatio-temporal patching to unify diverse data into a sequential format, facilitating the utilization of the powerful Transformer architecture. Thirdly, drawing inspiration from large language and vision models (Devlin et al., 2018; He et al., 2022), UniST adopts the widely-used generative pre-training strategy – Masked Token Modeling (MTM). We further enhance the model’s capability to capture complex spatio-temporal relationships by employing multiple masking strategies that comprehensively address multi-perspective correlations. Moreover, informed by the established domain knowledge in spatio-temporal modeling, we design an innovative prompt learning approach. The elaborated prompt network identifies underlying and shared spatio-temporal patterns, adapting dynamically to generate useful prompts. In this way, UniST aligns distinct data distributions of various datasets and advances towards developing a one-for-all universal model. We summarize our contributions as follows:
  • To our best knowledge, this the first attempt to address universal spatio-temporal prediction by investigating the potential of a one-for-all model in diverse spatio-temporal scenarios.
  • We propose UniST that harnesses data diversity and achieves universal spatio-temporal prediction through advanced pre-training and prompt learning. It has made a paradigm shift from traditional separate deep learning methods to a one-for-all model.
  • Extensive experiments demonstrate the generality and universality of UniST. It achieves new state-of-the-art performance on various prediction tasks, particularly, superior transferability in previously unseen scenarios.
@@ 2. Related Work
Urban spatio-temporal prediction (Wang et al., 2020; Zheng et al., 2014) aims to model and forecast the dynamic patterns of urban activities over space and time. Deep learning techniques has propelled significant advancements. A spectrum of models, including CNNs (Li et al., 2018; Zhang et al., 2017; Liu et al., 2018), RNNs (Wang et al., 2017, 2018; Lin et al., 2020), ResNets (Zhang et al., 2017), MLPs (Zhang et al., 2023b; Shao et al., 2022b), GNNs (Zhao et al., 2019; Bai et al., 2020; Geng et al., 2019), Transformers (Chen et al., 2022; Jiang et al., 2023; Yu et al., 2020; Chen et al., 2021), and diffusion models (Zhou et al., 2023a; Yuan et al., 2023), have been introduced to capture spatio-temporal patterns. Simultaneously, cutting-edge techniques like meta-learning (Lu et al., 2022; Yao et al., 2019), contrastive learning (Ji et al., 2023; Zhang et al., 2023a), and adversarial learning (Ouyang et al., 2023; Tang et al., 2022) are also utilized. However, most approaches remain constrained by training separate models for each specific dataset. Some studies (Yao et al., 2019; Lu et al., 2022; Jin et al., 2022; Liu et al., 2023d) explore transfer learning between cities, however, a certain amount of data samples in the target city are still required. Current solutions are restrictive to specified spatio-temporal scenarios and require training data, while our model allows generalization across diverse scenarios and provides a one-for-all solution.
Table 1. Comparison of UniST with other spatio-temporal models regarding important properties.
Model Scalability(1) Few-shot Zero-shot Efficicency
PromptST (Zhang et al., 2023c)
GPT-ST (Li et al., 2023b)
STEP (Shao et al., 2022a)
ST-SSL (Ji et al., 2023)
TrafficBERT (Jin et al., 2021)
TFM (Wang et al., 2023b)
UrbanGPT (Li et al., 2024) (2) (2)
STG-LLM (Liu et al., 2024)
UniST
  • (1)
    Whether can leverage diverse datasets with diverse formats.
  • (2)
    Restricted in the same city.
@@
Inspired by the remarkable strides in foundation models for NLP (Touvron et al., 2023; Brown et al., 2020) and CV (Rombach et al., 2022; Bai et al., 2023), foundation models for urban prediction have emerged recently. Some explorations unlock the potential of large language models (LLMs) in this context. Intelligent urban systems like CityGPT (Feng et al., 2024a; Xu et al., 2023), CityBench (Feng et al., 2024b) and UrbanGPT (Li et al., 2024) have demonstrated proficiency in addressing language-based tasks. Additionally, LLMs are utilized for describing urban-related images (Yan et al., 2023) to benefit downstream tasks and predict user activities (Gong et al., 2024). Moreover, the application of LLMs extends to traffic signal control (Lai et al., 2023), showcasing their utility in tackling complex spatio-temporal problems beyond languages. Recently, there also has been great progress in foundation models for time series (Jin et al., 2023b; Cao et al., 2023; Jin et al., 2023a; Zhou et al., 2023b). Unlike time series characterized by a straightforward sequential structure, spatio-temporal data presents a more intricate nature with intertwined dependencies across both spatial and temporal dimensions. While exploring the integration of LLMs is promising, it’s important to recognize that spatio-temporal data is not inherently generated by language. Thus, developing foundation models specifically trained on pure spatio-temporal data is also an important direction. In Table  1, we compare the essential properties of UniST with other approaches employing pre-training, prompt learning, or LLMs. UniST encompasses all these essential capabilities, whereas other approaches have certain limitations.
Prompt Learning
@@
Prompt learning has achieved superior performance in large models (Liu et al., 2023c; Qiao et al., 2022; Jia et al., 2022; Shen et al., 2024), with the goal of enhancing the generalization capability of pretrained models on specific tasks or domains. Typically, language models usually use a limited number of demonstrations as prompts and vision models often employ a learnable prompt network to generate useful prompts, known as prompt learning. Our research aligns with prompt learning, where spatio-temporal prompts are adaptively generated based on spatio-temporal patterns through a prompt network.
@@
Figure 2. The overview architecture of UniST, which consists of two stages: (i) large-scale spatio-temporal pre-trianing, (ii) spatio-temporal knowledge-guided prompt learning.
@@ 3. The UniST Model
3.1. Preliminary
Spatial and Temporal Partitions. We use a grid system for spatial partitioning, dividing the city into equal, non-overlapping areas defined by longitude and latitude on an $H\times W$ map. For each area, the temporal dynamics are recorded at certain intervals.
Spatio-Temporal Data. A spatio-temporal data $X$ is defined as a four-dimensional tensor with dimensions $T\times C\times H\times W$, where $T$ represents time steps, $C$ represents the number of variables, $H$ and $W$ represent spatial grids. $T$, $C$, $H$, and $W$ can vary across different spatio-temporal scenarios.
Spatio-Temporal Prediction. For a specific dataset, given $l_{h}$ historical observations for the grid map, we aim to predict the future $k$ steps. The spatio-temporal prediction task can be formulated as learning a $\theta$-parameterized model $\mathcal{F}$: $X_{[t:t+k]}=\mathcal{F}_{\theta}(X_{[t-l_{h}:t]})$.
Few-Shot and Zero-Shot Predictions. The model is trained on multiple source datasets and then adapted to a target dataset. In few-shot learning, it is fine-tuned with a small amount of target samples; in zero-shot learning, it makes predictions without any fine-tuning.
3.2. Pre-training and Prompt Learning
Universal spatio-temporal prediction aims to empower a single model to effectively handle diverse spatio-temporal scenarios, requiring the unification of varied spatio-temporal data within a cohesive model. This necessitates addressing significant distribution shifts across datasets of different scenarios. To achieve this goal, we propose a framework for pre-training and prompt learning, leading to a universal prediction model, UniST. Figure 2 shows the overview architecture, detailing UniST with two stages:
  • Stage 1: Large-scale spatio-temporal pre-training. Different from existing methods limited to a single dataset, our approach utilizing extensive spatio-temporal data from a variety of domains and cities for pre-training.
  • Stage 2: Spatio-temporal knowledge-guided prompt learning. We introduces a prompt network for in-context learning, where the generation of prompts is adaptively guided by well-developed spatio-temporal domain knowledge, such as spatial hierarchy and temporal periodicity.